# Finte-size induced random switching of chimeras in a deterministic two-population Kuramoto-Sakaguchi model

### v1

@@ Appendix A Justification for the approximations (24)
We present here a more formal justification for the approximations (24) used to close the mean-field dynamics (20). We begin by expressing
@@
For the synchronised population with $\alpha=1$ we have $r_{1}\approx 1$ which immediately yields the first approximation in (24). For $\alpha=2$, the bound (45) is less sharp.
@@
$z_{\alpha}^{(1,1)}$ $=\frac{1}{N}\sum_{j=1}^{N}\omega^{(\alpha)}_{j}e^{i\theta^{(\alpha)}_{j}}$
$=\langle\omega^{(\alpha)}_{j}\rangle_{j\in P_{\alpha}}\left(\frac{1}{N}\sum_{j=1}^{N}e^{i\theta^{(\alpha)}_{j}}\right)$
@@
Again for the synchronised population with $\alpha=1$ and $r_{1}\approx 1$ the second approximation in (24) follows immediately.
@@
for $\alpha=1,2$ and $m,n\in\mathbb{N}$, we can show that they remain satisfied if so at any time. We provide the motivation behind the assumptions (A) in Appendix A.1. Note that for $(m=0,n=0)$ and $(m=1,n=0)$ the conditions are trivially satisfied, and $(m=2,n=0)$ and $(m=1,n=1)$ correspond to (24).
Using the unperturbed dynamics (20) the conditions (A) imply that
@@
$=m\frac{1}{2}(v_{\alpha}z_{\alpha}^{m-1}-v_{\alpha}^{*}z_{\alpha}^{(2,0)}z_{\alpha}^{m-1}),$ (49)
for $\alpha=1,2,$ and $v_{\alpha}$ defined in (21). Since (A) implies $z_{\alpha}^{(m,0)}z_{\alpha}^{m^{\prime}}=z_{\alpha}^{(m+{m^{\prime}},0)}$ for all $m,m^{\prime}\in\mathbb{N}$, (A) simplifies to
$\frac{d}{dt}z_{\alpha}^{(m,n)}=m(iz_{\alpha}^{(m,n+1)}+\frac{1}{2}(v_{\alpha}z_{\alpha}^{(m-1,n)}-v_{\alpha}^{*}z_{\alpha}^{(m+1,n)})),$ (50)
for $n=0$ and $\alpha=1,2$. Note that (50) is trivially valid for $n\neq 0$, and corresponds to the unperturbed dynamics (23). This implies that if the conditions (A) hold true for any point in time, then they must be true for all times. We remark that there is an infinite family of conditions with this self-consistency property, which we describe in Appendix A.2.
@@ A.1 Motivation for conditions (A)
$\langle e^{im\theta^{(\alpha)}}\rangle$ $=(\langle e^{im\theta^{(\alpha)}}\rangle)^{m}.$ (52)
@@
$=\frac{\int_{-\pi}^{\pi}\frac{e^{im\theta}}{\nu^{\alpha}}d\theta}{\int_{-\pi}^{\pi}\frac{1}{\nu^{(\alpha)}}d\theta}$
$=\int_{-\pi}^{\pi}\frac{e^{im\theta}}{2\pi}\frac{\sqrt{(\bar{\omega}^{(\alpha)}-\Omega)^{2}-|v_{\alpha}|^{2}}}{\bar{\omega}^{(\alpha)}-\Omega+\text{Im}(v_{\alpha}e^{-i\theta})}d\theta$ (53)
@@
$p$ $=\frac{\bar{\omega}^{(\alpha)}-\Omega}{\left|v_{\alpha}\right|}-\sqrt{\frac{\bar{\omega}^{(\alpha)}-\Omega}{\left|v_{\alpha}\right|}-1}\sqrt{\frac{\bar{\omega}^{(\alpha)}-\Omega}{\left|v_{\alpha}\right|}+1},$
@@
$\langle e^{im\theta^{(\alpha)}}\rangle$ $=\int_{-\pi}^{\pi}\frac{e^{im\theta}}{2\pi}\frac{1-p^{2}}{1+p^{2}-2p\operatorname{Re}\left({\rm{v}}e^{-i\theta}\right)}d\theta$
$=\int_{-\pi}^{\pi}\frac{e^{im\theta}}{2\pi}\frac{1-p^{2}}{1+p^{2}-p\left(\frac{{\rm{v}}}{e^{i\theta}}+\frac{e^{i\theta}}{{\rm{v}}}\right)}d\theta$
$=\int_{-\pi}^{\pi}\frac{e^{im\theta}}{2\pi}\left(1+\frac{p\frac{{\rm{v}}}{e^{i\theta}}}{1-p\frac{{\rm{v}}}{e^{i\theta}}}+\frac{p\frac{e^{i\theta}}{{\rm{v}}}}{1-p\frac{e^{i\theta}}{{\rm{v}}}}\right)d\theta$
$=\int_{-\pi}^{\pi}\frac{e^{im\theta}}{2\pi}\sum_{n=-\infty}^{\infty}p^{\left|n\right|}\left(\frac{{\rm{v}}}{e^{i\theta}}\right)^{n}d\theta$
@@
With $\bar{\omega}=0$ the approximations (A) immediately follow.
@@ A.2 Self-consistent approximations for $z^{(m,n)}_{\alpha}$
Here we derive a more general set of approximations for $z^{(m,n)}_{\alpha}$ which satisfies the set of self-consistency conditions (A). We restrict the approximations to be of the form
@@
where $f_{m}$ is an unknown function to be determined. Special cases are $f_{0}(r_{\alpha})=1$ and $f_{1}(r_{\alpha})=r_{\alpha}$.
@@
Additionally, we remark that (60) holding true for $m=2$, and (59) holding true for all $m$, implies that (60) holds true for all $m$. This means that to find a valid set of functions $f_{m}$, we must only find some $f_{2}$ that satisfies
@@
$f_{2}(r_{\alpha})=2r_{\alpha}\frac{r_{\alpha}(4-g^{2})+g\sqrt{4-r_{\alpha}^{2}(4-g^{2})}}{4}-1,$ (62)
@@
Note that for $c=0$ and $g=\infty$ (A.2) yields condition (A) with $z_{\alpha}^{(m,n)}=0^{n}z_{\alpha}^{m}$.
@@ Appendix B Numerical integrator for the OU process
Consider a general $k$-dimensional stochastic OU process
@@
where $\Gamma\in\mathbb{R}^{k\times k}$ and $\Sigma\in\mathbb{R}^{k\times k}$ are the drift and diffusion matrices, respectively, and $W_{t}$ is a $k$-dimensional Wiener process.
@@
with $\Delta W\sim\mathcal{N}(0,\Delta t\,\operatorname{I})$ and
@@
$\hat{\Sigma}$ $=\sqrt{\frac{\int_{0}^{\Delta t}\exp(-\Gamma s)\Sigma\Sigma^{\top}\exp(-\Gamma^{\top}s)\,ds}{\Delta t}}.$ (68)
@@
The integral over the matrix exponential in (68),
$\int_{0}^{t}\exp(-\Gamma s)\Sigma\Sigma^{\top}\exp(-\Gamma^{\top}s)\,ds,$ (69)
can be explicitly calculated for diagonalizable drift matrices $\Gamma=VDV^{-1}$ with diagonal $D$ as
@@
In the special case where drift and diffusion matrices are in the form (33)-(34), (70) simplifies to
@@
We remark that for more general cases the analytical formula for the variance can be costly to compute in high dimensions, and one may be well advised to employ a naive approach of repeated Euler-Maruyama steps Sachs, Leimkuhler, and Danos (2017).
@@ Appendix C Additional numerical results
To show that our stochastic modelling framework is generally valid, we present results for a different choice of parameters with $K=85$, $\kappa=55$ and $\lambda=\pi/2-0.09$. For these parameters we obtain $\langle Z\rangle=0.689+0.131\,i$ from the mean-field theory. Fitting the covariance function by minimising the cost function (35) from $t_{\rm{min}}=0.5$ to $t_{\rm{max}}=1$ yields $\gamma=1.918-18.05i$ and $\sigma=0.335$.
@@
We show in Figures 1923 that our stochastically driven model (V)–(30) reproduces the statistics of the order parameter and the phases very well. The case of a small network with $N=8$ is less well approximated. Figure 23 shows the mean rotation frequency ${\bar{\nu}}_{i}$ for various values of $N$. We used the mean-field result for $\langle Z\rangle$ in the reduced stochastic model (V)–(30). We remark that if we used the empirical mean estimated from simulations of the two-population KS model (1)–(2) for each value of $N$ the mean rotation frequencies are as accurate as those shown in Figure 17. Figure 24 shows that the azimuthally averaged equation for the fluctuations is capable to provide good estimates of the mean switching time using the Kramers problem solution (43).
@@
Figure 24: Comparison of the mean switching time $\bar{\tau}$ for different system sizes NN, obtained from a long-time simulations of the full two-population KS model (1)–(2) (blue circles) and from the analytical expression (40) and (43) based on solving Kramers problem using our reduced stochastic model. We show the result of the Kramers problem for $q^{\star}=0.261$ (orange line) which was estimated from the full two-population KS model (1)–(2), and for $q^{\star}=0.236$ (yellow line) which was estimated from the reduced deterministic mean-field model (IV). Equation parameters are as in Figure 19.

### v3

@@ Appendix A Justification for the approximations (26)
We present here a more formal justification for the approximations (26) used to close the mean-field dynamics (IV). We begin by expressing
@@
For the synchronised population with $\alpha=1$, we have $r_{1}\approx 1$, which immediately yields the first approximation in (26). For $\alpha=2$, the bound (47) is less sharp.
@@
$z_{\alpha}^{(1,1)}$ $=\frac{1}{N}\sum_{j\in P_{\alpha}}\omega^{(\alpha)}_{j}e^{i\theta^{(\alpha)}_{j}}$
$=\langle\omega^{(\alpha)}_{j}\rangle_{j\in P_{\alpha}}\left(\frac{1}{N}\sum_{j\in P_{\alpha}}e^{i\theta^{(\alpha)}_{j}}\right)$
@@
For the synchronised population with $\alpha=1$ and $r_{1}\approx 1$ the second approximation in (26) then follows immediately.
@@
for $\alpha=1,2$ and $m,n\in\mathbb{N}$, we can show that they remain satisfied if so at any time. We provide the motivation behind the assumptions (A) in Appendix A.1 below. Note that for $(m=0,n=0)$ and $(m=1,n=0)$ the conditions are trivially satisfied, and $(m=2,n=0)$ and $(m=1,n=1)$ correspond to (26).
Using the unperturbed dynamics (IV), conditions (A) imply that
@@
$=m\frac{1}{2}(v_{\alpha}z_{\alpha}^{m-1}-v_{\alpha}^{*}z_{\alpha}^{(2,0)}z_{\alpha}^{m-1})$ (51)
for $\alpha=1,2,$ and $v_{\alpha}$ defined in (9)–(10). Since (A) implies $z_{\alpha}^{(m,0)}z_{\alpha}^{m^{\prime}}=z_{\alpha}^{(m+{m^{\prime}},0)}$ for all $m,m^{\prime}\in\mathbb{N}$, (A) simplifies to
$\frac{d}{dt}z_{\alpha}^{(m,n)}=m(iz_{\alpha}^{(m,n+1)}+\frac{1}{2}(v_{\alpha}z_{\alpha}^{(m-1,n)}-v_{\alpha}^{*}z_{\alpha}^{(m+1,n)}))$ (52)
for $n=0$ and $\alpha=1,2$. Note that (52) is trivially valid for $n\neq 0$, and corresponds to the unperturbed dynamics (25). This implies that if conditions (A) hold for any point in time, then they must be true for all times. We remark that there is an infinite family of conditions with this self-consistency property, which we describe in Appendix A.2 below.
@@ A.1 Motivation for conditions (A)
$\langle e^{im\theta^{(\alpha)}}\rangle$ $=(\langle e^{i\theta^{(\alpha)}}\rangle)^{m}.$ (54)
@@
$=\frac{1}{\int_{-\pi}^{\pi}\frac{1}{\nu^{(\alpha)}}d\theta}\,\int_{-\pi}^{\pi}\frac{e^{im\theta}}{\nu^{\alpha}}d\theta$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{im\theta}\frac{\sqrt{(\bar{\omega}^{(\alpha)}-\Omega)^{2}-|v_{\alpha}|^{2}}}{\bar{\omega}^{(\alpha)}-\Omega+\text{Im}(v_{\alpha}e^{-i\theta})}d\theta.$ (55)
@@
$p$ $=\frac{\bar{\omega}^{(\alpha)}-\Omega}{\left|v_{\alpha}\right|}-\sqrt{\frac{\bar{\omega}^{(\alpha)}-\Omega}{\left|v_{\alpha}\right|}-1}\sqrt{\frac{\bar{\omega}^{(\alpha)}-\Omega}{\left|v_{\alpha}\right|}+1},$ (56)
@@
$\langle e^{im\theta^{(\alpha)}}\rangle$ $=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{im\theta}\frac{1-p^{2}}{1+p^{2}-2p\operatorname{Re}\left({\rm{v}}e^{-i\theta}\right)}d\theta$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{im\theta}\frac{1-p^{2}}{1+p^{2}-p\left(\frac{{\rm{v}}}{e^{i\theta}}+\frac{e^{i\theta}}{{\rm{v}}}\right)}d\theta$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{im\theta}\left(1+\frac{p\frac{{\rm{v}}}{e^{i\theta}}}{1-p\frac{{\rm{v}}}{e^{i\theta}}}+\frac{p\frac{e^{i\theta}}{{\rm{v}}}}{1-p\frac{e^{i\theta}}{{\rm{v}}}}\right)d\theta$
$=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{im\theta}\sum_{n=-\infty}^{\infty}p^{\left|n\right|}\left(\frac{{\rm{v}}}{e^{i\theta}}\right)^{n}d\theta$
@@
With $\bar{\omega}=0$ conditions (A) immediately follow.
@@ A.2 Self-consistent approximations for $z^{(m,n)}_{\alpha}$
Here we derive a more general set of approximations for $z^{(m,n)}_{\alpha}$, which satisfies the set of self-consistency conditions (A). We restrict the approximations to be of the form
@@
where $f_{m}$ is an unknown function to be determined. Special cases are $f_{0}(r_{\alpha})=1$ and $f_{1}(r_{\alpha})=r_{\alpha}$. The temporal evolution of $z_{\alpha}^{(m,n)}$ is given by
@@
Note that if (63) holds for $m=2$ and (62) holds for all $m$, then (63) holds for all $m$. This means that to find a valid set of functions $f_{m}$, we must only find some $f_{2}$ that satisfies
@@
$f_{2}(r_{\alpha})=2r_{\alpha}\frac{r_{\alpha}(4-g^{2})+g\sqrt{4-r_{\alpha}^{2}(4-g^{2})}}{4}-1$ (65)
@@
Note that for $c=0$ and $g=\infty$ (A.2) yields conditions (A) with $z_{\alpha}^{(m,n)}=0^{n}z_{\alpha}^{m}$.
@@ Appendix B Numerical integrator for the OU process
Consider a general $d$-dimensional stochastic OU process
@@
where $\Gamma\in\mathbb{R}^{d\times d}$ and $\Sigma\in\mathbb{R}^{d\times d}$ are the drift and diffusion matrices, respectively, and $W_{t}$ is a $d$-dimensional Wiener process.
@@
with $\Delta W\sim\mathcal{N}(0,\Delta t\operatorname{I}_{d})$ and
@@
$\hat{\Sigma}$ $=\frac{1}{\sqrt{\Delta t}}\sqrt{\int_{0}^{\Delta t}\exp(-\Gamma s)\Sigma\Sigma^{\top}\exp(-\Gamma^{\top}s)\,ds}.$ (71)
@@
The integral over the matrix exponential in (71) can be explicitly calculated for diagonalizable drift matrices $\Gamma=VDV^{-1}$ with diagonal $D$ as
@@
For the special case of drift and diffusion matrices that are of the form (35)–(36), (72) simplifies to
@@
We remark that for more general cases, the analytical formula for the variance can be costly to compute in high dimensions, and one may be well advised to employ a naive approach of repeated Euler-Maruyama steps Sachs, Leimkuhler, and Danos (2017).
@@ Appendix C Additional numerical results
To show that our stochastic modelling framework is generally valid, we present results for a different choice of parameters with $K=85$, $\kappa=55$ and $\lambda=\pi/2-0.09$. For these parameters we obtain $\langle Z\rangle=0.689+0.131\,i$ from the mean-field theory. Fitting the covariance function by minimising the cost function (37) in the range from $t_{\rm{min}}=0.5$ to $t_{\rm{max}}=1$ yields $\gamma=1.918-18.05i$ and $\sigma=0.335$.
@@
We show in Figures 1822 that our stochastically driven model (V)–(32) reproduces the statistics of the order parameter and the phases very well. The case of a small network with $N=8$ is less well approximated. Figure 22 shows the mean rotation frequency ${\bar{\nu}}_{i}$ for various values of $N$. We used the mean-field result for $\langle Z\rangle$ in the reduced stochastic model (V)–(32). We remark that if we used the empirical mean estimated from simulations of the two-population KS model (1)–(2) for each value of $N$, the mean rotation frequencies are as accurate as those shown in Figure 16 in the main manuscript. Figure 23 shows that the azimuthally averaged equation for the fluctuations is capable to provide good estimates of the mean switching time using the Kramers problem solution (45).
Appendix D Random draws of the native frequencies
We present results for the case where the native frequencies are drawn randomly rather than equiprobably, i.e. where the two populations are not exactly the same. We show results for the same set of parameters as used in the main manuscript.
The empirical histograms of the order parameters are not very sensitive to the particular realisation of the native frequencies. We show in Figure 24 that our stochastically driven model (V)–(32) (using the same fitted values of $\gamma=1.75-25.34\,i$ and $\sigma=0.574$ as used in the main manuscript) reproduces the statistics of the order parameter and the phases very well even in the case of random samplings of native frequencies $\omega_{i}$. As expected, the case of a small network with $N=8$ is less well approximated.
Similarly, on the level of reproducing the statistics of individual oscillators, Figures 2527 show that the stochastically driven model (V)–(32) reproduces the mean and the variance of the fluctuations of individual synchronised oscillators around their mean values as well as their mean rotation frequencies.
Whereas the statistics of the order parameter and the fluctuations are not very sensitive to the particular realisation of the native frequencies, the probability of switching events sensitively depends on the particular realisation of the draw. In Figure 28 we show time series of the order parameters $r_{1,2}$ for $N=10$ oscillators for two different realisations of the native frequencies drawn from $\omega\sim{\mathcal{N}}(0,1)$. Whereas for one realisation we observe frequent random switching during the $1,000$ time units, another realisation exhibits no switching events during the same period. This illustrates that the probability of switching to occur in a given time window crucially depends on the particular realisation of the native frequencies. To explain this, we hypothesise that the variance of a particular native frequency draw is a major factor determining this probability. More concretely, we hypothesise that a larger spread of the native frequencies increases the tail in the distribution of the fluctuations $\zeta$, responsible for inducing switching. Conversely, we hypothesise that realisations with significantly smaller variance in one population increases the stability of its synchronised state, and hence decreases the probability of random switches. Figure 29 shows the observed number of switches for $400$ random native frequency draws, according to their variances ${\rm{Var}}(\omega^{(1)})$ and ${\rm{Var}}(\omega^{(2)})$. It is clearly seen that realisations that do not exhibit any switching in the given time period are characterised by native frequency draws where at least one of the populations has a significantly smaller than expected variance (denoted by purple points). Moreover, the figure supports our hypothesis that increased overall variance promotes switching. The average switching frequency across all $400$ realisations, $\langle\bar{\tau}^{-1}\rangle_{\omega\sim g(\omega)}=1/137.6$, is approximately equal to the empirical switching frequency $\bar{\tau}^{-1}=133.0^{-1}$ for equiprobably drawn natural frequencies as reported in the main manuscript. This supports the conclusion that equiprobable samples provide a representative realisation of the distribution $g(\omega)$. Note that we report here on the frequency of switching rather than on the switching times, because of the occurrence of realisations which do not exhibit any switching in the given time window.
@@ Appendix D Random draws of the native frequencies
Figure 22: Mean rotation frequency ${\bar{\nu}}_{i}$ of the synchronised oscillators as a function of their native frequencies $\omega_{i}$ for the full two-population KS model (1)–(2) (blue circles) and for the reduced stochastic model (V)–(32) (red crosses) for $N=8,12,16$ and $32$. The dashed lines denote the mean frequency $\Omega$ as calculated from the self-consistency relations (19)–(20) of the mean-field theory. Equation parameters are as in Figure 18.
Figure 23: Comparison of the mean switching time $\bar{\tau}$ for different system sizes NN, obtained from a long-time simulations of the full two-population KS model (1)–(2) (blue circles) and from the analytical expressions (42) and (45) based on solving Kramers problem using our reduced stochastic model. We show the result of the Kramers problem for $q^{\star}=0.261$ (red line) which was estimated from the full two-population KS model (1)–(2), and for $q^{\star}=0.236$ (yellow line) which was estimated from the reduced deterministic mean-field model (IV). Equation parameters are as in Figure 18.
Figure 24: Comparison of the empirical histograms for the order parameter $r_{1}$ of the synchronised population obtained from a single trajectory of the full two-population KS model (1)–(2) (blue histogram) and from the reduced stochastic model (V)–(32) (red line) for $N=8,12,16$ and $32$, where the native frequencies $\omega_{i}$ are drawn randomly and independently from a standard normal distribution $\mathcal{N}(0,1)$. Equation parameters are $K=100$, $\kappa=60$ and $\lambda=\frac{\pi}{2}-0.075$.
Figure 25: Mean of the fluctuations of the synchronised phases around their mean, $\theta_{i}-\psi_{1}$, as a function of their native frequencies $\omega_{i}$ for the full two-population KS model (1)–(2) (blue circles) and for the reduced stochastic model (V)–(32) (red crosses) for $N=8,12,16$ and $32$, where the native frequencies $\omega_{i}$ are drawn randomly and independently from a standard normal distribution $\mathcal{N}(0,1)$. Equation parameters are as in Figure 1 of the main manuscript.
Figure 26: Variance of the fluctuations of the synchronised phases around their mean, $\theta_{i}-\psi_{1}$, as a function of their native frequencies $\omega_{i}$ for the full two-population KS model (1)–(2) (blue circles) and for the reduced stochastic model (V)–(32) (red crosses) for $N=8,12,16$ and $32$, where the native frequencies $\omega_{i}$ are drawn randomly and independently from a standard normal distribution $\mathcal{N}(0,1)$. Equation parameters are as in Figure 1 of the main manuscript.
Figure 27: Mean rotation frequency ${\bar{\nu}}_{i}$ of the synchronised oscillators as a function of their native frequencies $\omega_{i}$ for the full two-population KS model (1)–(2) (blue circles) and for the reduced stochastic model (V)–(32) (red crosses) for $N=8,12,16$ and $32$, where the native frequencies $\omega_{i}$ are drawn randomly and independently from a standard normal distribution $\mathcal{N}(0,1)$. The dashed lines denote the mean frequency $\Omega$ as calculated from the self-consistency relations (19)–(20) of the mean-field theory. Equation parameters are as in Figure 1 of the main manuscript.
Figure 28: Order parameters $r_{1}$ (blue) and $r_{2}$ (red) for the two-population KS model (1)–(2) as a function of time for $N=10$ oscillators. We show results for two different realisations of native frequencies drawn randomly and independently from a normal distribution with unit variance, $g(\omega)\sim{\mathcal{N}}(0,1)$. Equation parameters are as in Figure 1 of the main manuscript.
Figure 29: Dependency of the probability of chimera switching on the realisation of the native frequency draw. Each of $400$ independent realisations was integrated for $2,000$ time units and the number of switching events in that time were recorded. Realisations, characterised by the variances of their respective populations, ${\rm{Var}}(\omega^{(1)})$ and ${\rm{Var}}(\omega^{(2)})$, are coloured according to the number of observed switching events; purple denotes realisations for which no switching has occurred. Equation parameters are as in Figure 1 of the main manuscript.